Calculate the buoyant force on a cube of metal with an edge of 1.3 cm that is placed in salt water. The density of the metal is 7.86 x 103 kg/m², and the density of the salt water is 1.025 x 103 kg/


1
Expert's answer
2021-01-29T13:22:55-0500

By the definition of the buoyant force, we have:


"F_B=\rho_wgV=\rho_wga^3,""F_B=1.025\cdot10^3\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot(0.013\ m)^3=0.022\ N."